OpenStudy (anonymous):

One gallon of fuel mixture contains antifreeze in the ratio of 5 parts fuel to one part antifreeze. To this is added half a gallon of mixture which is one third antifreeze and two thirds fuel. What is the ratio of fuel to antifreeze in the final mixture? (fuel/antifreeze)

OpenStudy (hba):

first, since you are mixing one gallon with a half gallon, change the full gallon (5:1 mixture) to 2 half gallons, each (5:1). Now you have 3 separate half gallons: 2 with 5/6 fuel, 1/6 antifreeze, and 1 with 4/6 fuel and 2/6 antifreeze. I converted each measurements to sixths for easy adding. Next, just add all of the parts together: half gallon 1 + half gallon 2 + half gallon 3 = 5/6 fuel + 1/6 antifreeze + 5/6 fuel + 1/6 antifreeze + 4/6 fuel + 2/6 antifreeze =14/6 fuel + 4/6 antifreeze 14/6 + 4/6 = 18/6 = 3 "half gallons" ratio is 14:4 = 7:2

OpenStudy (anonymous):

OHHH gosh, that was easy! Thanks for clearing that up!!!

OpenStudy (anonymous):

@hba ^^

OpenStudy (hba):

@Areesha.1D You are quiet welcome :)

OpenStudy (anonymous):

@hba *quite

OpenStudy (hba):

@Areesha.1D Grammar Nazi

OpenStudy (anonymous):

@hba your label is 100% correct ^_^

OpenStudy (hba):

@Areesha.1D I had to add a word before it but i can't :)

OpenStudy (hba):

@Areesha.1D Its can't* not couldn't

OpenStudy (hba):

@Areesha.1D Wait i will pm you :D